OpenStudy (anonymous):

Nathan is trying out to be the kicker on his high school's football team. He notices that for every hour he practices he can kick the ball 15 feet further. On his first try he kicked the ball 30 feet, on the second try he kicked the ball 15 yards. What could be a possible distance for his third kick? A. 15 feet B. 30 feet C. 20 yards D. 30 yards

OpenStudy (anonymous):

c

OpenStudy (anonymous):

15 yards is 45 feet and 20 yards is 60 feet

OpenStudy (anonymous):

The cost, in dollars, of an order of deli meat is a function of its weight w in pounds. The function c(w) outputs the cost c of an order of meat weighing w pounds. How can the statement c(1.5) = 9 be interpreted? A. The cost c is $6.00 for 9 pounds of meat. B. The cost of 9 pounds of meat is $1.50. C. The cost of 1.5 pounds of meat is $9.00. D. The cost c is $6.00 for 1.5 pounds of meat.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Because he kicks it 15 feet farther every time.

OpenStudy (anonymous):

Ariana has $4 to spend at the candy shop. She wants at least 10 sour snaps which are 15 cents each, and at least 3 chocolate truffles which are 80 cents each. Which of the following systems of inequalities correctly represents the constraints on the variables in this problem? Let s represent the sour snaps and t represent the chocolate truffles. A. B. C. D.

OpenStudy (anonymous):

Can't see answers.

OpenStudy (anonymous):

Yeah links wont work

OpenStudy (anonymous):

Henry receives a flat rate of $100 per day and can earn a $2 commission for each computer repair (r) he performs. Yolanda receives a flat rate of $80 per day and can earn $5 per computer sale (s) she makes. How would the two equations model the wages Henry and Yolanda make each day? A. H=100-2r Y=80-5s B. H=80+2r Y=100+5s C. H=100+2r Y=80+5s D. H=2r-100 Y=5s+80

OpenStudy (anonymous):

Had to type answers lol

OpenStudy (anonymous):

C.